# Fail Fast, Win Big: Rethinking the Drafting Strategy in Speculative Decoding via Diffusion LLMs

### v2

@@ 5.1 Setup
Models and Baselines. In our main experiments, we evaluate on three target models of different sizes: Qwen2.5-{32B,14B,7B}-Instruct (Yang et al., 2024). The different speculative decoding schemes we evaluate are listed below. For all schemes, we set the generation temperature of all models to 0 and set the maximum number of output tokens to 1024. We discuss more details of our best-effort reproduction of all baselines and additional baselines in Appendix D.
@@
Fast-dLLM (Wu et al., 2025a): For the dLLM drafter baseline, we employ Fast_dLLM_v2_1.5B (Wu et al., 2025a), the state-of-the-art diffusion LLM that embeds efficiency optimizations such as approximate caching and confidence-aware parallel decoding. We adopt its default hyperparameters that achieve the best balance between accuracy and generation speed. This baseline represents naively plugging in a dLLM as the draft model without the additional optimizations of FailFast.
EAGLE-3 (Li et al., 2025): EAGLE-3 is one of the most widely deployed state-of-the-art speculative decoding drafters. As a lightweight single-layer autoregressive Transformer, EAGLE-3 introduces a multi-layer feature fusion and a training-time test mechanism for improving drafting quality. Similar to our other baselines, we also pick the best speculation length for each model-dataset combination.
FailFast: For the hyperparameters in Alg. 1, we set for the {32,14,7}B target model and . Although picking specific hyperparameters for each dataset/model combination yields higher speedups, we use the same set of hyperparameters across datasets to demonstrate generalizability. FailFast’s performance degrades gracefully as hyperparameters are tweaked (Appendix C).
@@ 5.2 End-to-End Results
Workload sensitivity. Across models and datasets, FailFast achieves an overall speedup of 1.9-4.9. Compared to baselines, FailFast achieves a speedup of 1.3-2.1 over the AR drafter baseline and a speedup of 1.2-1.7 over the Fast-dLLM baseline. FailFast has a higher win over baselines on datasets that are easier to speculate: Math and coding workloads exhibit centralized regions of predictable tokens (e.g., intermediate calculations and code generation), allowing FailFast to capitalize on long speculation windows. Conversely, datasets where “easy” tokens are scattered rather than clustered offer fewer opportunities for aggressive speculation length expansion, though FailFast still outperforms fixed-length baselines from a moderate level of length expansion.
@@
Efficiency of FailFast. Naively plugging in a dLLM as the drafter (Fast-dLLM) retains the acceptance rate (Tab. 3) and achieves a speedup over the AR drafter. However, as shown in Fig. 4, this gain is primarily derived from a reduction in speculation latency (22.9% on average)111This improvement is less drastic than the theoretical end-to-end speedups dLLMs can offer (up to 2.5 reported) (Wu et al., 2025a) because our termination criterion requires the next continuous tokens to be unmasked; due to the random decoding order within token blocks, the leftmost tokens are not necessarily decoded first. while the verification latency remains near-constant.
@@
In contrast, FailFast further reduces speculation latency significantly (a further 41.1% over Fast-dLLM) by employing an ultra-small compute budget for the drafter. Although our approach inevitably degrades the acceptance rate (Tab. 3) – a combined result of low drafter compute budget and proposing more tokens on average in each round – we achieve superior end-to-end speedups (e.g., 3.1-4.9 for Qwen2.5-32B). In addition, FailFast reduces verification latency (Fig. 4) by an average of 17.1% via adopting adaptive speculation lengths: through dynamically expanding the speculation window in “easy” regions, we can speculate and verify up to 70 tokens (Fig. 6) in a single round. This drastic reduction in the total number of speculation-verification rounds and the speculation overhead in each round outweighs the cost of a lower per-token acceptance rate. We include an example trajectory in Appendix F to visualize FailFast in action.
@@
Detailed Analyses. Due to space constraints, we only further dissect FailFast’s performance improvements on Qwen2.5-32B on the MATH dataset, and we report the data on all other datasets and model combinations in Appendix B. Even though the AR Drafter has a higher overall acceptance rate (Tab. 3), for each of its forward passes, it produces only one token and accepts 0.6 tokens on average. In contrast, FailFast can produce an average of 4.1 tokens per forward pass and accept 1.7 of them. Another source of FailFast’s win comes from its aggressive speculation length expansion. Fig. 6 shows a CDF of the accepted/speculated length across rounds. In 20% of the rounds, we extended the speculation length beyond the default, reducing the average number of speculation-verification rounds by 16.0%.
@@
Comparisons with EAGLE-3. Across models, FailFast achieves an average speedup of 1.2-1.7 over EAGLE-3 (up to 1.4-2.0 on specific datasets). Even though EAGLE-3 has a smaller drafting latency in each of its forward passes – it only has one layer compared to the 28 layers in other baselines, reducing the kernel launch overhead – it remains an autoregressive drafter and only generates a single token in each forward pass, whereas FailFast generates multiple tokens in each forward pass. Further, EAGLE-3 uses a fixed speculation length and is not able to capitalize on regions of easier tokens as a result. Even if a token tree with, for example, 64 tokens was used in EAGLE-3 drafting, the upper bound on accepted tokens would still be bottlenecked by the depth of the tree (up to 8 in EAGLE-3 (Li et al., 2025)).
@@
Due to space constraints, we refer the readers to the Appendix sections for a more in-depth discussion regarding EAGLE-3, the implementation details of baselines, a hyperparameter sensitivity analysis of FailFast, and a discussion of explorations of alternative design choices.
@@ 6 Conclusion
Acknowledgments We thank Princeton’s Systems for Artificial Intelligence Lab (SAIL) and Princeton Language and Intelligence (PLI) for providing the hardware resources for running experiments. Rui would like to thank the many full-time researchers and interns at Google SRG for their help and support, especially Nikos Pagonas, Howie Chen, Manos Chatzakis, Vinay Banakar, Liana Patel, Jiani Zhang, and Lily Tsai. In addition, Rui would like to thank Minghao Yan for a random conversation on speculative decoding back in 2024 that planted the seed of this work.
@@ Appendix A List of abbreviations and notations
| / | Speculation length (num. tokens) in each round |
@@ Appendix C Hyperparameter Sensitivity Analysis of FailFast
Maximum Speculation Length (). We observe that values between 30 and 70 perform reliably well. We adopt a default value of because LLM prefill for short sequences is typically memory-bandwidth bound, and the typical tipping point between memory-bound and compute-bound for medium-sized LLMs typically ranges between 64 and 128 tokens on modern GPUs (Agrawal et al., 2024). In high-throughput inference settings where batch sizes are large, bigger values of still work, although they might inflate the verification latency slightly – the prefill pass would eventually become compute-bound for large sequence lengths, even with prefix caching (Pan et al., 2025b). In that case, synchronous verification might force a reduction in to not inflate the per-round verification latency, and FailFast’s performance improvement from “winning big” might drop slightly. However, FailFast’s performance improvement from “failing fast” persists. We note that speculative decoding is inherently more suitable for inference acceleration at smaller batch sizes to begin with and provides diminishing gains with an increase in batch size, where target model inference becomes less memory-bound and more compute-bound (Sadhukhan et al., 2024; amd, 2025).
@@ D.3 EAGLE-3
All EAGLE-3 baselines are run using vLLM v0.13.0 (Kwon et al., 2023). We run an extensive parameter sweep of num_speculative_tokens from 3 to 20, and pick a different speculation length that yielded the best speedup for each model-dataset pair. There are no public EAGLE-3 weights available for Qwen2.5. Thus, we conducted pretraining of EAGLE-3 weights using the ultrachat-200K (Ding et al., 2023) text corpus in SpecForge (spe, 2025). We adopted the default training hyperparameters in SpecForge and trained EAGLE-3 to match the target model’s output until convergence. We plot the moving average of the training accuracy in Fig. 7. During evaluation, we use the model checkpoints after five epochs of training (260k training steps). We find that even though further training improves training-time accuracy, they have a negligible impact on the end-to-end speedup of EAGLE-3. We open-source our EAGLE-3 model weights and training scripts.
Draft token trees. The vLLM EAGLE-3 implementation only supports greedy decoding, where a single sequence of draft tokens is submitted to the target model for verification. To ensure a fair and accurate representation of EAGLE-3’s peak performance (Li et al., 2025; eag, 2025), we also attempted to run the same EAGLE-3 weights in SGLang v0.5.6.post2 (Zheng et al., 2023) with a draft token tree of 64 tokens (to roughly match in FailFast). However, the performance is not as strong as that for vLLM. We hypothesize that the underlying reason is that even though EAGLE-3 speculates a total of 64 tokens in each tree, the depth of the tree is not as deep compared to FailFast – EAGLE-3 uses a default draft tree depth of 8 (Li et al., 2025) – and as a result, EAGLE-3 was not able to capitalize from regions of easier tokens, and the additional tokens in its draft tree led to undue speculation latency. Blindly adopting a fixed speculation length of 64 tokens in vLLM resulted in a significant slowdown compared to vanilla decoding. We note that FailFast’s one-step generation is compatible with token trees, which can be realized via a custom attention mask (Agrawal et al., 2025). We leave this extension as future work.
@@
Data-matched comparisons of parametric drafters. We acknowledge that the training data used for our dLLM drafter and our EAGLE baseline do not exactly match. Our dLLM drafter was adapted from Qwen2.5-1.5B-Instruct via a block-wise diffusion training process (Wu et al., 2025a). As a result, the difference in their performance might be partially attributed to this factor (although the comparisons with AR-drafters are indeed data-matched). Due to the significant amount of resources needed for a rigorous controlled experiment of different drafting methods, where training needs to be data-matched, parameter-matched, and FLOP-matched (Waleffe et al., 2024; Dao & Gu, 2024), we leave it as future work and use this paper as a first step in demonstrating dLLM drafter’s potential.

### v3

@@ 5.1 Setup
Models and Baselines. In our main experiments, we evaluate three target models of different sizes: Qwen2.5-{32B,14B,7B}-Instruct (Yang et al., 2024). The different speculative decoding schemes we evaluate are listed below. For all schemes, we set the generation temperature of all models to 0 and set the maximum number of output tokens to 1024. We discuss more details of our best-effort reproduction of all baselines and additional baselines in Appendix D.
@@
Fast-dLLM (Wu et al., 2025a): For the dLLM drafter baseline, we employ Fast_dLLM_v2_1.5B (Wu et al., 2025a), the state-of-the-art semi-autoregressive diffusion LLM that embeds efficiency optimizations such as approximate caching and confidence-aware parallel decoding. We adopt its default hyperparameters that achieve the best balance between accuracy and generation speed. This baseline represents naively plugging in a dLLM as the draft model without the additional optimizations of FailFast.
EAGLE-3 (Li et al., 2025): EAGLE-3 is one of the most widely deployed state-of-the-art speculative decoding drafters. As a lightweight single-layer autoregressive Transformer, EAGLE-3 introduces a multi-layer feature fusion and a training-time test mechanism for improving drafting quality. Different from other strategies where a single chain of draft tokens is proposed for verification, EAGLE-3 leverages a draft-tree-based mechanism that improves the overall token acceptance rate.
FailFast: FailFast employs Fast_dLLM_v2_1.5B as the underlying model and performs one-step generation and dynamic speculation length expansion. For the hyperparameters in Alg. 1, we set for the {32,14,7}B target model and . Although picking specific hyperparameters for each dataset/model combination yields higher speedups, we use the same set of hyperparameters across datasets to demonstrate generalizability. FailFast’s performance degrades gracefully as hyperparameters are tweaked (Appendix C).
@@ 5.2 End-to-End Results
Workload sensitivity. Across models and datasets, FailFast achieves an overall speedup of 1.9-4.9. Compared to baselines, FailFast achieves a speedup of 1.3-2.1 over the AR drafter baseline and a speedup of 1.2-1.7 over the Fast-dLLM baseline. FailFast has a higher win over baselines on datasets that are easier to speculate: math and coding workloads exhibit centralized regions of predictable tokens (e.g., intermediate calculations and code generation), allowing FailFast to capitalize on long speculation windows. Conversely, datasets where “easy” tokens are scattered rather than clustered offer fewer opportunities for aggressive speculation length expansion, though FailFast still outperforms fixed-length baselines with moderate length expansion.
@@
Efficiency of FailFast. Naively plugging in a dLLM as the drafter (Fast-dLLM) retains the acceptance rate (Tab. 5) and achieves a speedup over the AR drafter. However, as shown in Fig. 6, this gain is primarily derived from a reduction in speculation latency (22.9% on average)333This improvement falls short of the theoretical end-to-end speedups of dLLMs (Wu et al., 2025a) because our termination criterion requires the next continuous tokens to be unmasked; due to the random decoding order within token blocks, the leftmost tokens are not necessarily decoded first. while the verification latency remains near-constant.
@@
In contrast, FailFast further reduces speculation latency significantly (a further 41.1% over Fast-dLLM) by employing an ultra-small compute budget for the drafter. Although our approach inevitably degrades the acceptance rate (Tab. 5) – a combined result of a low drafter compute budget and proposing more tokens on average in each round – we achieve superior end-to-end speedups (e.g., 3.1-4.9 for Qwen2.5-32B). In addition, FailFast reduces verification latency (Fig. 6) by an average of 17.1% via adopting adaptive speculation lengths: through dynamically expanding the speculation window in “easy” regions, we can speculate and verify up to 70 tokens (Fig. 8) in a single round. This drastic reduction in the total number of speculation-verification rounds and the speculation overhead in each round outweighs the cost of a lower per-token acceptance rate. We include an example trajectory in Appendix F to visualize FailFast in action.
@@
Detailed analyses. Due to space constraints, we only further dissect FailFast’s performance improvements on Qwen2.5-32B on the MATH dataset, and we report the data on all other datasets and model combinations in Appendix B. Even though the AR drafter has a higher overall acceptance rate (Tab. 5), for each of its forward passes, it produces only one token and accepts 0.6 tokens on average. In contrast, FailFast can produce an average of 4.1 tokens per forward pass and accept 1.7 of them. Another source of FailFast’s win comes from its aggressive speculation length expansion. Fig. 8 shows a CDF of the accepted/speculated length across rounds. In 20% of the rounds, we extended the speculation length beyond the default, reducing the average number of speculation-verification rounds by 16.0%. Interestingly, FailFast’s dynamic policy converges to an average speculation length of 14.6 tokens, closely matching the theoretically optimal of 16 tokens from our prior analysis (Fig. 2).
@@
Comparisons with EAGLE-3. Across models, FailFast achieves an average speedup of 1.0–1.4 over EAGLE-3, with gains of up to 1.3–1.7 on specific datasets. Even though EAGLE-3 incurs a smaller drafting latency in each of its forward passes – it only has one layer compared to the 28 layers in other drafters, reducing overheads associated with loading weights and launching kernels – FailFast still outperforms EAGLE-3. This is because EAGLE-3 remains an autoregressive drafter that generates only one token per forward pass, whereas FailFast generates multiple tokens in each pass. Moreover, while EAGLE-3 speculates a tree of 50–60 tokens per round, its draft tree is relatively shallow (default depth 8 (Li et al., 2025)) compared to FailFast, which drafts a chain of up to 70 tokens in a single round. As a result, EAGLE-3’s upper bound on the number of accepted tokens is constrained by tree depth, limiting its ability to exploit extended regions of easier tokens.
@@
Due to space constraints, we refer readers to the Appendix sections for a more in-depth discussion regarding EAGLE-3, the implementation details of baselines, a hyperparameter sensitivity analysis of FailFast, and a discussion of explorations of alternative design choices.
@@ 6 Conclusion
Impact Statement This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgments We thank Princeton’s Systems for Artificial Intelligence Lab (SAIL) and Princeton Language and Intelligence (PLI) for providing the hardware resources for running experiments. Rui would like to thank the many full-time researchers and interns at Google SRG for their help and support, especially Nikos Pagonas, Howie Chen, Manos Chatzakis, Vinay Banakar, Liana Patel, Jiani Zhang, and Lily Tsai. In addition, Rui would like to thank Minghao Yan for a random conversation on speculative decoding back in 2024 that planted the seed of this work and for providing feedback on an earlier version of this paper.
@@ Appendix A List of abbreviations and notations
| / | Speculation length (number of tokens) in each round |
@@ Appendix C Hyperparameter Sensitivity Analysis of FailFast
Maximum Speculation Length (). We observe that values between 30 and 70 perform reliably well. We adopt a default value of because LLM prefill for short sequences is typically memory-bandwidth bound, and the typical tipping point between memory-bound and compute-bound for medium-sized LLMs typically ranges between 64 and 128 tokens on modern GPUs (Agrawal et al., 2024). In high-throughput inference settings where batch sizes are large, bigger values of still work, although they might inflate the verification latency slightly – the prefill pass would eventually become compute-bound for large sequence lengths, even with prefix caching (Pan et al., 2025b). In that case, synchronous verification might force a reduction in to avoid inflating the per-round verification latency, and FailFast’s performance improvement from “winning big” might drop slightly. However, FailFast’s performance improvement from “failing fast” persists. We note that speculative decoding is inherently more suitable for inference acceleration at smaller batch sizes to begin with and provides diminishing gains as batch size increases, where target model inference becomes less memory-bound and more compute-bound (Sadhukhan et al., 2024; amd, 2025).
@@ D.3 EAGLE-3
EAGLE-3 training. There are no public EAGLE-3 weights available for Qwen2.5. Thus, we conducted pretraining of EAGLE-3 weights using the ultrachat-200K (Ding et al., 2023) text corpus in SpecForge (spe, 2025). We adopted the default training hyperparameters in SpecForge and trained EAGLE-3 to match the target model’s output until convergence. We plot the moving average of the training accuracy in Fig. 9. During evaluation, we use the model checkpoints after five epochs of training (260k training steps). We find that even though further training improves training-time accuracy, they have a negligible impact on the end-to-end speedup of EAGLE-3. We open-source our EAGLE-3 model weights and training scripts.
Draft token trees. To ensure a fair and accurate representation of EAGLE-3’s peak performance (Li et al., 2025; eag, 2025), we evaluate EAGLE-3 in SGLang v0.5.6.post2 (Zheng et al., 2023) with a draft token tree, with 50-60 tokens in each tree (following the best default hyperparameter configurations reported in the EAGLE-3 paper (Li et al., 2025), roughly matching in FailFast). For completeness, we additionally evaluate EAGLE-3 configured to propose a single chain of 3-20 draft tokens (rather than a tree). This variant consistently underperforms the full draft-tree version of EAGLE-3, and we therefore omit its results from the evaluation. Similar to the AR drafter (Fig. 2), adopting a fixed speculation length of 50-60 tokens in a single chain of draft tokens resulted in a significant slowdown compared to vanilla decoding due to the speculation overhead in autoregressive drafting. We note that FailFast’s one-step generation is compatible with token trees, which can be realized via a custom attention mask (Agrawal et al., 2025). We leave this extension as future work.
@@
Data-matched comparisons of parametric drafters. We acknowledge that the training data used for our dLLM drafter and our EAGLE baseline do not exactly match. Our dLLM drafter, Fast_dLLM_v2_1.5B, was adapted from Qwen2.5-1.5B-Instruct via a block-wise diffusion training process (Wu et al., 2025a). As a result, the difference in their performance might be partially attributed to this factor (although the comparisons with AR-drafters are indeed data-matched). Due to the significant amount of resources needed for a rigorous controlled experiment of different drafting methods, where training needs to be data-matched, parameter-matched, and FLOP-matched (Waleffe et al., 2024; Dao & Gu, 2024), we leave it as future work and use this paper as a first step in demonstrating dLLM drafter’s potential.